s. To the nearest tenth of a mis, what is the final velocity of the car in the previous problem? 7. A passenger in a helicopter initially traveling upwards at 26 m/s accidentally drops a package out the window. If it takes 19 seconds to reach the ground (y-0), how high to the nearest meter was the helicopter when the package was dropped? To the nearest meter what was the maximum height of the package above the ground in the previous problem? 8. 9. A speeder traveling at 42 m/s passes a motorcycle policeman at rest a of the road. The second how long does it take the policeman to catch the speeder? policeman accelerates at 3.23 m/s. To the nearest tenth of a 10. How far to the nearest tenth of a meter can a runner running at 12 ms run in the time it takes a rock to fall from rest 56 meters?
Answer #1

7.

Using 2nd equation of motion

h = ho + ut - 0.5 gt^2

0 = ho + 26 *19 - 0.5*9.8* 19^2

ho = 1274.9 m

8.

ho = 1274.9 + 26^2/2 x 9.8

ho = 1309.39 m

Comment in case any doubt please rate my answer .....

According to the HOMEWORKLIB RULES I can answer one question at a time so please post other questions seperately....